6. Later we will give data on an experiment conducted by Mosteller and Wallace (1963) on the use of language by Alexander Hamilton. Their study attempted to unravel a longstanding question of who wrote the 12 disputed Federalist Papers. We will show that if X- the number of times Hamilton used the word an in a given text, then a good distribution function for that random variable is 0, ifx<0 F(x)-〈 y, (1.24)%-1.24 Using this result, compute the following probabilities. (a) PrXso
Answer #1

a)P(X<=0)=P(X=0)=(1.24)0e-1.24/.0! =e-1.24 =0.289384

b)P(0<X<1) =0 (As there is not discrete value in between)

c)P(0<X<=1)=P(X=1)=(1.24)1e-1.24/.1! =0.358836

d)P(0<=X<1)=P(X=0)=(1.24)0e-1.24/.0! =e-1.24 =0.289384